A basketball player is standing on the floor 10.0
Answer #1

a) At any point in a freefalling objects trajectory, the acceleration is 9.81 m/s² downwards

b)

The horizontal position of the ball (from the player) is (V0*cos37.6º)*t, where t is the travel time. To travel 10.0 m, the time is then t = 10.0/(V0*cos37.6º) = 12.62/V0

The ball must reach the height of the basket in that time, starting at height of 2.18 m. The height of the ball is h = h0 + V0*t*sin38.0º - ½*g*t²

3.05 - 2.18 = 0.610*12.62 - ½*9.8*159.26/V0²
6.828= 780.3/V0²
V0 = 10.68 m/s